OpenStudy (anonymous):

Solve: 4(x + 6) = 6(x − 2)

OpenStudy (anonymous):

show your work

OpenStudy (anonymous):

Distributive property of multiplication: \[a(b+c) = ab + ac\]

OpenStudy (anonymous):

please

OpenStudy (anonymous):

can you get rid of the parentheses in both sides now?

OpenStudy (anonymous):

no may you please show me it would be very much appreacitated

OpenStudy (anonymous):

a(b+c)=ab+ac

OpenStudy (anonymous):

So write the equation for me without the parentheses, it will get easier from there

OpenStudy (anonymous):

4(x+6) = ?

OpenStudy (anonymous):

are you there?